OpenStudy (anonymous):

a^2+2b=7 b^2+4c=-7 c^2+ab=-14 then find a^2 + b^2 + c^2

OpenStudy (shamil98):

a^2 + 2b = 7 a^2 = 7 - 2b b^2 + 4c = -7 b^2 = -4c-7 c^2 + ab = -14 c^2 = -ab - 14 go from there

OpenStudy (anonymous):

but the answer is given as 14 ..numerical .!

OpenStudy (nikato):

what lesson is this based on?

OpenStudy (anonymous):

the chapter is equations ..:/

OpenStudy (agent0smith):

Probably have to use substitution to solve. Looks gross.

OpenStudy (anonymous):

there are second degree terms too ...substitution will be a lot tiring ..! okay if substituion then how ..approach ?

OpenStudy (shamil98):

good luck...

OpenStudy (anonymous):

i cant do that ..! so should i close it then

OpenStudy (agent0smith):

I agree with @shamil98 ...olving three linear equations is bad enough. I'm sure it's doable with substitution. I am not willing to spend time trying to show you how to do it.

OpenStudy (anonymous):

even me too ..:/ okay .!

OpenStudy (haseeb96):

 this is the answer